# Spin-up and spin distribution of stellar black holes grown by gas accretion in proto-stellar clusters

### 19 Mar 2026 v1

@@ 3 BH-spin evolution and accretion disk
Our primary goal is to quantify the possible spin-up of stellar BHs through accretion in proto-stellar clusters. A stellar BH is formed with low spin from its progenitor star if there is efficient angular-momentum transfer. The magnetic Tayler instability may efficiently couple the stellar core to its envelope (Spruit, 2002), enabling angular momentum to be transported outward during core contraction and thereby preventing the core from spinning up. Simulations indicate that the BH formed at collapse is expected to have a very low natal spin, with a typical value of $a_{*}\simeq 0.01$ (Fuller and Ma, 2019), where the dimensionless spin is $a_{*}\equiv J_{\rm BH}c/(Gm_{\rm BH}^{2})$ and $J_{\rm BH}$ the BH spin. We adopt this value for the initial BH spin magnitudes.
Regarding the initial BH spin orientations, we assume an isotropic distribution inherited from an isotropic distribution of stellar spin axes. This assumption is consistent with observational analyses (Jackson and Jeffries, 2010; Healy et al., 2023). It is also intuitively justified, at least for non-rotating clusters, by the turbulent nature of star-forming environments. We note, nevertheless, that rotating clusters with sufficient rotational energy may induce partially stellar spin alignment (Corsaro et al., 2017). We do not consider this case here.
@@ 3.1 Disk formation
The BH motion is repeatedly perturbed at stochastic timesteps, discussed in Appendix (A). The acquired BH azimuthal velocity induces apparent transverse velocity gradients within the accretion sphere ($|\bm{r}-\bm{r}_{\bullet}|\leq R_{\rm acc}$) as the sphere gets advected along the BH trajectory. Numerical simulations of angular-momentum accretion, driven by velocity and density gradients, show the formation of an accretion disk (Ruffert and Anzer, 1995). More recently, Tripathi et al. (2025) demonstrated that an off-axis (“lateral”) Bondi-Hoyle-Lyttleton flow, characterized by an antisymmetric transverse velocity field, naturally generates angular momentum and produces a persistent disk.
@@
The velocity shear across $R_{\rm acc}$ produces a net angular momentum as the apparent gas transverse velocity in the BH rest frame differs between the inner (proximate to the cluster center) and outer hemispheres, creating a shear flow opposite to the BH’s orbital motion. This asymmetric inflow provides angular-momentum injection within the BH’s accretion radius $R_{\rm acc}$. At the BH orbital plane this shear specific angular momentum of the gas is $\bm{j}_{0}\approx-2\bm{\omega}_{\rm BH}R_{\rm acc}^{2}$. Volume-averaging over all planes within $R_{\rm acc}$ (using $R_{\rm plane}=R_{\rm acc}\cos\theta$, $\theta$ measured from the orbital plane) and weighting by the volume element $\propto\cos^{2}\theta$, we obtain the geometric prefactor: $2\int_{-\pi/2}^{\pi/2}\cos^{4}\theta d\theta/\int_{-\pi/2}^{\pi/2}\cos^{2}\theta d\theta=3/2$. Thus, we get the total specific shear angular momentum injected into the accretion radius
@@
where the minus sign indicates that the shear generates retrograde inflow relative to BH orbital motion. The accretion radius is defined in (7). Both the orientation and approximate magnitude of our estimate, equation (10), is consistent with numerical simulations of angular-momentum accretion driven by velocity gradients (Ruffert and Anzer, 1995; Tripathi et al., 2025).
Note that all BH-dependent quantities are local and time-dependent, e.g. $\bm{j}=\bm{j}(\bm{r}_{\bullet},t)$, although we omit this dependence to simplify our notation. There is also an angular momentum component ($\bm{j}_{\rm grad}$) induced by the density gradient within $R_{\rm acc}$. For the smooth cluster density profiles in our models, this is negligible because $\bm{j}_{\rm grad}\approx(\Delta m_{\rm in-out}/m_{\rm acc})\bm{\omega}_{\rm BH}R_{\rm acc}^{2}$ with $\Delta m_{\rm in-out}/m_{\rm acc}\ll 1$ (mass difference between inner and outer accretion hemispheres over total sphere mass). It is straightforward to verify that for our typical parameters of a Plummer sphere, a BH inside the Plummer radius and a disk of any radius $R_{\rm d}<10^{-3}\,{\rm pc}$ we get $j_{\rm grad}/j\lesssim 10^{-3}$. We therefore neglected the $j_{\rm grad}$ contribution to improve computational efficiency.
In our system, a rotationally supported disk can only form if the gas circularizes inside the BH’s radius of influence but outside the innermost stable circular orbit (ISCO)
@@
We define the circularization radius as
@@ 3.2 Spin magnitude evolution
In our system, the accretion rate
@@
depends sensitively not only on BH mass, but also on BH position and velocity. Therefore, it varies for the same BH during the evolution of the system. We have
$0\leq\dot{m}(t)\leq 10,$ (15)
for each BH. A geometrically thin-disk approximation is valid for stellar BHs if $\dot{m}\lesssim 0.3$ (Reynolds, 2021). For higher accretion rates, the disk is mildly geometrically thick or “slim” up to our upper accretion cap $\dot{m}\leq 10$ (Sądowski, 2009). Thus, we assume that the disk of each BH is
@@
and use the appropriate model at each instant of time.
A slim disk has lower radiative efficiency $\eta$ than a thin disk (Sądowski, 2009). We adopt the simple prescription (Inayoshi et al., 2016)
@@
This gives $\eta_{\rm eff}\approx\eta_{\rm thin}$ at $\dot{m}\lesssim 0.3$. At the cap $\dot{m}=10$ we get $\eta_{\rm eff}=0.014-0.08$ ($\eta_{\rm thin}=0.057-0.32$) for spin zero and one, respectively.
We calculate the BH mass increase rate as
$\dot{m}_{\rm BH}=(1-\eta_{\rm eff}(a_{*}))\dot{m}_{\rm acc}.$ (17)
and the BH-spin magnitude evolution
@@
$\Lambda\equiv\left\{\begin{array}[]{ll}\Lambda_{\rm thin},&\dot{m}<0.3\\ \Lambda_{\rm slim},&\dot{m}\geq 0.3\end{array}\right$. (19)
The values of $\Lambda_{\rm thin}\equiv\Lambda(R_{\rm ISCO})$, $\eta_{\rm thin}\equiv 1-\mathcal{E}(R_{\rm ISCO})$ are given by the standard thin-disk expressions (Bardeen et al., 1972; Thorne, 1974).
We introduce a phenomenological relation to estimate the effective $\Lambda_{\rm slim}$ for slim disks
@@
where $f_{\Lambda}(H/R)$ decreases linearly. This linear dependence is motivated by general relativistic magnetohydrodynamic (GRMHD) simulations (Penna et al., 2010), which show a linear decrease of the specific angular momentum accreted by the BH with increasing disk thickness $H/R$ of a few percent. We estimate $H/R$, modelling the simulation data of the slim-disk models of Sądowski et al. (2011a) with a simple saturating prescription
@@
This gives reasonable values $H/R=\{0.07,\,0.12,\,0.25\}$ for $\dot{m}=\{0.3,\,1,\,10\}$, respectively. We have $f_{\Lambda}^{\rm max}=1$ for the thin disk $\dot{m}=0.3$. Instead, the parameter $f_{\Lambda}^{\rm min}$ is defining the minimum possible angular momentum inflow corresponding to our maximum allowed accretion rate $\dot{m}=10$. The direct correspondence between $f_{\Lambda}$ and $a_{*,{\rm max}}$, which can be derived by the equilibrium condition of (18), allows for the estimation of $f_{\Lambda}^{\rm min}$ by estimating $a_{*,\rm{max}}$ for slim disks $\dot{m}=10$. The analytical calculations of Sądowski et al. (2011b) suggest that for slim disks at super-Eddington accretion rates it is $a_{*,\rm{max}}\approx 0.99$ close to the Thorne limit. Instead, GRMHD simulations of slim disks show that we may get a significantly lower equilibrium spin $a_{*,\rm{max}}\approx 0.93$ (Gammie et al., 2004). This corresponds to $f_{\Lambda}^{\rm min}=0.95$. We adopt this value as a conservative assumption on the spin-up process and perform sensitivity checks. We find that our results are robust and practically unaffected for values even down to $f_{\Lambda}^{\rm min}=0.88$, corresponding to the extreme case of an equilibrium spin $a_{*,\rm{max}}=0.90$ at $\dot{m}=10$ for every single accretion cycle at this rate. Even for $f_{\Lambda}^{\rm min}=0.80$ our derived spin-mass correlation, including the high-spin plateau, remains robust, while the saturation spin drops only by $\lesssim 4\%$.
@@
We note that if a magnetically arrested disk (MAD) (Bisnovatyi-Kogan and Ruzmaikin, 1974) forms, the accreting BH spins down due to angular momentum extraction via the Blandford-Znajek mechanism (Tchekhovskoy et al., 2011). Ricarte et al. (2023) have estimated that for thick disks the equilibrium spin drops down to $a_{*,{\rm max}}\approx 0.8,\,0.2$ for $\dot{m}=1,\,10$, respectively. However, in the dense stochastic environment of a gaseous proto-cluster the maintenance of a long-lived coherent MAD is uncertain. In addition, Ricarte et al. (2023) suggest that the spin equilibration timescale for MAD accretion attains the values $t_{\rm eq}^{\rm MAD}\approx 45,\,10,\,4.5\,{\rm Myr}$ for $\dot{m}=1,\,5,\,10$, respectively. Our mass-gap BHs have been accreting during their evolution at a rate $\dot{m}\sim 1-3$ or less, while only the IMBH with $m_{\rm BH}\sim 10^{3}\,{\rm M}_{\odot}$ may have reached up to $\dot{m}\sim 10$. In our model the cluster gas is depleted within $13\,{\rm Myr}$, therefore MAD accretion is certainly irrelevant for our mass-gap BHs ($150\,{\rm M}_{\odot}$) for which $t_{\rm eq}^{\rm MAD}>t_{\rm depletion}$. Most importantly, as we discuss in Appendix A, the BHs are disrupted at a timescale $\Delta t_{\rm stoch}\approx 10^{-2}\,{\rm Myr}$. Therefore, both for mass-gap and IMBHs, even if a MAD disk forms, it cannot remain coherent after a few disruption events and does not have sufficient time during its lifetime to cause any significant spin-down. Overall, our model assumes that MAD accretion is negligible.
@@ 3.3 Spin-disk orientation
The accretion disk, if formed, gets erratically disrupted, repeatedly, every timestep $\Delta t_{\rm stoch}$ (dynamically calculated as the system evolves), when a stochastick kick due to stellar encounters and gas turbulence is applied. Since the disk orientation depends on the BH motion, each disruption event drives a disk re-orientation.
In order to evaluate the spin-disk relative orientation after every disruption event, we need the disk angular momentum $J_{\rm d}\equiv J(R_{\rm d})$ where
@@
Therefore, we need a model for the surface density $\Sigma$ and the disk radius $R_{\rm d}$ which does not necessarily coincide with the circularization radius. We calculate the disk surface density adopting the Shakura-Sunjaev model (Shakura and Sunyaev, 1973; Frank et al., 2002)
@@
where $f(R)=(1-(R/R_{\rm ISCO})^{-1/2})^{1/4}$ and the gravitational radius is $r_{\rm g}=Gm_{\rm BH}/c^{2}$. For thin disks we adopt the viscosity parameter $\alpha_{\rm thin}=0.1$. For slim disks, the Shakura-Sunjaev surface density can be a good approximation using an effective viscosity parameter $\alpha_{\rm eff}=0.01$ (Sądowski et al., 2011a). Indeed, we verify that for $\alpha=0.01$, the Shakura-Sunyaev surface density matches the slim-disk GRRMHD simulations of Sądowski et al. (2011a) to within a factor of order unity. Therefore we adopt this as a sufficient analytic approximation. We also perform exploratory sensitivity checks over the range $\alpha_{\rm eff}=0.001-0.1$. The general trend of our derived spin-mass correlation is robust. A notable result is that higher $\alpha_{\rm eff}$ produce a somewhat broader BH spin distribution for the more massive BHs. Especially for $\alpha_{\rm eff}=0.1$ the IMBHs ($\sim 10^{3}{\rm M}_{\odot}$) exhibit reduced lower and upper spin boundaries and a nearly uniform spin distribution within this range, $a_{*}\sim 0.3-0.8$. For $m_{\rm BH}\lesssim 800\,{\rm M}_{\odot}$, however, there is no reduction in spin boundaries, but only a slight increase in their spread. Nevertheless, we emphasize that $\alpha_{\rm eff}=0.1$ is not strictly self-consistent within our slim-disk treatment, because it lowers the effective optical depth, thereby shifting the validity of the model below our super-Eddington regime for slim disks (Sądowski et al., 2011a).
@@
In order to define the disk radius $R_{\rm d}$, we need first to introduce three characteristic radii. Firstly, we define the outer radius $R_{\rm out}$ as the maximum outer radius at which torques have enough time to be communicated, before the BH motion gets disrupted by a stochastic kick. It is defined by the condition $t_{\rm w}(R_{\rm out})=\Delta t_{\rm stoch}$ which gives
@@
$\frac{R_{\rm out}}{r_{\rm g}}=\left\{\begin{array}[]{ll}1.5\cdot 10^{8}\left(\frac{\Delta t_{\rm stoch}}{10^{-3}\,{\rm Myr}}\right)^{\frac{4}{5}}\left(\frac{m_{\rm BH}}{50{\rm M}_{\odot}}\right)^{-\frac{24}{25}}&,\;\mbox{thin}\\ 3.5\cdot 10^{8}\left(\frac{\Delta t_{\rm stoch}}{10^{-3}\,{\rm Myr}}\right)^{\frac{2}{3}}\left(\frac{H/R}{0.1}\right)^{\frac{2}{3}}\left(\frac{m_{\rm BH}}{50{\rm M}_{\odot}}\right)^{-\frac{2}{3}}&,\;\mbox{slim}\end{array}\right$. (25)
@@
We define $t_{\rm w}$ and discuss timescales in Appendix B. It should be $R_{\rm d}\leq R_{\rm out}$, otherwise there is not enough time for the torque to be communicated out to $R_{\rm d}$ within the stochastic timestep $\Delta t_{\rm stoch}$.
@@
The Toomre radius is defined by the condition $Q(R_{\rm Q})=1$, where the Toomre parameter is $Q(R)=c_{\rm s,disk}\Omega_{\rm K,disk}(R)/(\pi G\Sigma(R))$. Stable disks are constrained by the condition $Q(R)>1$ that is $R<R_{\rm Q}$. We have
@@
$\frac{R_{\rm Q}}{r_{\rm g}}=\left\{\begin{array}[]{ll}1.7\cdot 10^{10}\left(\frac{m_{\rm BH}}{50{\rm M}_{\odot}}\right)^{-\frac{52}{45}}\dot{m}^{-\frac{22}{45}}&,\;\mbox{thin}\\ 2.9\cdot 10^{9}\left(\frac{H/R}{0.1}\right)^{\frac{4}{5}}\left(\frac{m_{\rm BH}}{50{\rm M}_{\odot}}\right)^{-\frac{24}{25}}\dot{m}^{-\frac{14}{25}}&,\;\mbox{slim}\end{array}\right$. (26)
It should also be $R_{\rm d}\leq R_{\rm Q}$, otherwise the disk is unstable.
@@
At last, we define the “breaking” radius $R_{\rm break}$. Numerical evidence suggests that the disk (either thin or slim) breaks into rings above a threshold radius (Nixon et al., 2012; Nealon et al., 2015)
@@
$\frac{R_{\rm break}}{r_{\rm g}}=\left(\frac{4a_{*}}{3\alpha(H/R)}\left|\sin\theta_{\rm rel}\right|\right)^{2/3}\,,\quad 45^{o}\lesssim\theta_{\rm rel}\lesssim 135^{o}$ (27)
@@
for such severe misalignment (Nixon et al., 2012). The angle $\theta_{\rm rel}$ is the relative angle between the initial (at the time a stochastic kick is applied) disk angular momentum and BH-spin orientation
$\cos\theta_{\rm rel}=-\hat{\bm{\omega}}_{\rm BH}\cdot\hat{\bm{J}}_{BH}|_{t=t_{\rm ini}}$ (28)
@@
since $\hat{\bm{J}}_{\rm disk}=-\hat{\bm{\omega}}_{\rm BH}$. It should be $R_{\rm d}\leq R_{\rm break}$ when $45^{o}\lesssim\theta_{\rm rel}\lesssim 135^{o}$, otherwise the disk breaks into rings.
@@
Thus, $R_{\rm d}$ should be smaller than all four characteristic radii $R_{\rm circ}$, $R_{\rm out}$, $R_{\rm Q}$, $R_{\rm break}$, that is we define
$R_{\rm d}=\min\{R_{\rm circ},R_{\rm out},R_{\rm Q},R_{\rm break}\}$ (29)
@@
provided always that $R_{\rm ISCO}<R_{\rm circ}\leq R_{\rm acc}$ so that a disk is able to form. The $R_{\rm break}$ is considered only when $45^{o}\lesssim\theta_{\rm rel}\lesssim 135^{o}$; otherwise it is ignored.
@@
For $R_{\rm acc}=R_{\rm B}$ which is the typical case for BHs close the cluster center, the circularization radius is

### 29 Apr 2026 v2

@@ 3 BH-spin evolution and accretion disk
Our primary goal is to quantify the possible spin-up of stellar BHs through accretion in proto-stellar clusters. A stellar BH is formed with low spin from its progenitor star if there is efficient angular-momentum transfer. The magnetic Tayler instability may efficiently couple the stellar core to its envelope (Spruit, 2002), enabling angular momentum to be transported outward during core contraction and thereby preventing the core from spinning up. Simulations indicate that the BH formed at collapse is expected to have a very low natal spin, with a typical value of $a_{*}\simeq 0.01$ (Fuller and Ma, 2019), where the dimensionless spin is $a_{*}\equiv J_{\rm BH}c/(Gm_{\rm BH}^{2})$ and $J_{\rm BH}$ the BH spin. We adopted this value for the initial BH spin magnitudes.
Regarding the initial BH spin orientations, we assumed an isotropic distribution inherited from an isotropic distribution of stellar spin axes. This assumption is consistent with observational analyses (Jackson and Jeffries, 2010; Healy et al., 2023). It is also intuitively justified, at least for nonrotating clusters, by the turbulent nature of star-forming environments. We note, nevertheless, that rotating clusters with sufficient rotational energy may induce partially stellar spin alignment (Corsaro et al., 2017). We do not consider this case here.
@@ 3.1 Disk formation
The BH motion is repeatedly perturbed at stochastic timesteps, as is discussed in Appendix A. The acquired BH azimuthal velocity induces apparent transverse velocity gradients within the accretion sphere ($|\bm{r}-\bm{r}_{\bullet}|\leq R_{\rm acc}$) as the sphere gets advected along the BH trajectory. Numerical simulations of angular-momentum accretion, driven by velocity and density gradients, show the formation of an accretion disk (Ruffert and Anzer, 1995). More recently, Tripathi et al. (2025) demonstrated that an off-axis (“lateral”) Bondi-Hoyle-Lyttleton flow, characterized by an antisymmetric transverse velocity field, naturally generates angular momentum and produces a persistent disk.
@@
The velocity shear across $R_{\rm acc}$ produces a net angular momentum as the apparent gas transverse velocity in the BH rest frame differs between the inner (proximate to the cluster center) and outer hemispheres, creating a shear flow opposite to the BH’s orbital motion. This asymmetric inflow provides angular-momentum injection within the BH’s accretion radius $R_{\rm acc}$. At the BH orbital plane, this shear specific angular momentum of the gas is $\bm{j}_{0}\approx-2\bm{\omega}_{\rm BH}R_{\rm acc}^{2}$. Volume-averaging over all planes within $R_{\rm acc}$ (using $R_{\rm plane}=R_{\rm acc}\cos\theta$, $\theta$ measured from the orbital plane) and weighting by the volume element $\propto\cos^{2}\theta$, we obtain the geometric prefactor: $2\int_{-\pi/2}^{\pi/2}\cos^{4}\theta d\theta/\int_{-\pi/2}^{\pi/2}\cos^{2}\theta d\theta=3/2$. Thus, we get the total specific shear angular momentum injected into the accretion radius,
@@
where the minus sign indicates that the shear generates retrograde inflow relative to BH orbital motion. The accretion radius is defined in Eq. (7). Both the orientation and approximate magnitude of our estimate, Eq. (10), is consistent with numerical simulations of angular-momentum accretion driven by velocity gradients (Ruffert and Anzer, 1995; Tripathi et al., 2025).
Note that all BH-dependent quantities are local and time-dependent, such as $\bm{j}=\bm{j}(\bm{r}_{\bullet},t)$, although we omit this dependence to simplify our notation. There is also an angular momentum component ($\bm{j}_{\rm grad}$) induced by the density gradient within $R_{\rm acc}$. For the smooth cluster density profiles in our models, this is negligible because $\bm{j}_{\rm grad}\approx(\Delta m_{\rm in-out}/m_{\rm acc})\bm{\omega}_{\rm BH}R_{\rm acc}^{2}$ with $\Delta m_{\rm in-out}/m_{\rm acc}\ll 1$ (the mass difference between the inner and outer accretion hemispheres over the total sphere mass). It is straightforward to verify that for our typical parameters of a Plummer sphere, a BH inside the Plummer radius and a disk of any radius, $R_{\rm d}<10^{-3}\,{\rm pc}$, we get $j_{\rm grad}/j\lesssim 10^{-3}$. We therefore neglected the $j_{\rm grad}$ contribution to improve computational efficiency.
In our system, a rotationally supported disk can only form if the gas circularizes inside the BH’s radius of influence but outside the innermost stable circular orbit (ISCO),
@@
We defined the circularization radius as
@@ 3.2 Spin magnitude evolution
In our system, the accretion rate,
@@
depends sensitively not only on the BH mass, but also on the BH position and velocity. Therefore, it varies for the same BH during the evolution of the system. We have
$0\leq\dot{m}(t)\leq 10$ (15)
for each BH. A geometrically thin-disk approximation is valid for stellar BHs if $\dot{m}\lesssim 0.3$ (Reynolds, 2021). For higher accretion rates, the disk is mildly geometrically thick or “slim” up to our upper accretion cap, $\dot{m}\leq 10$ (Sądowski, 2009). Thus, we assumed that the disk of each BH is
@@
and used the appropriate model at each instant of time.
A slim disk has a lower radiative efficiency, $\eta$, than a thin disk (Sądowski, 2009). We adopted the simple prescription (Inayoshi et al., 2016)
@@
This gives $\eta_{\rm eff}\approx\eta_{\rm thin}$ at $\dot{m}\lesssim 0.3$. At the cap $\dot{m}=10$, we get $\eta_{\rm eff}=0.014-0.08$ ($\eta_{\rm thin}=0.057-0.32$) for spin zero and one, respectively.
We calculated the BH mass increase rate as
$\dot{m}_{\rm BH}=(1-\eta_{\rm eff}(a_{*}))\dot{m}_{\rm acc}$ (17)
and the BH-spin magnitude evolution as
@@
$\Lambda\equiv\left\{\begin{array}[]{ll}\Lambda_{\rm thin},&\dot{m}<0.3\\ \Lambda_{\rm slim},&\dot{m}\geq 0.3\end{array}\right.$. (19)
The values of $\Lambda_{\rm thin}\equiv\Lambda(R_{\rm ISCO})$ and $\eta_{\rm thin}\equiv 1-\mathcal{E}(R_{\rm ISCO})$ are given by the standard thin-disk expressions (Bardeen et al., 1972; Thorne, 1974).
We introduced a phenomenological relation to estimate the effective $\Lambda_{\rm slim}$ for slim disks,
@@
where $f_{\Lambda}(H/R)$ decreases linearly. This linear dependence is motivated by general relativistic magnetohydrodynamic (GRMHD) simulations (Penna et al., 2010), which show a linear decrease in the specific angular momentum accreted by the BH with increasing disk thickness, $H/R$, of a few percent. We estimated $H/R$, modeling the simulation data of the slim-disk models of Sądowski et al. (2011a) with a simple saturating prescription,
@@
This gives reasonable values of $H/R=\{0.07,\,0.12,\,0.25\}$ for $\dot{m}=\{0.3,\,1,\,10\}$, respectively. We have $f_{\Lambda}^{\rm max}=1$ for the thin disk, $\dot{m}=0.3$. Instead, the parameter $f_{\Lambda}^{\rm min}$ is defining the minimum possible angular momentum inflow corresponding to our maximum allowed accretion rate, $\dot{m}=10$. The direct correspondence between $f_{\Lambda}$ and $a_{*,{\rm max}}$, which can be derived by the equilibrium condition of (18), allows for the estimation of $f_{\Lambda}^{\rm min}$ by estimating $a_{*,\rm{max}}$ for slim disks, $\dot{m}=10$. The analytical calculations of Sądowski et al. (2011b) suggest that for slim disks at super-Eddington accretion rates it is $a_{*,\rm{max}}\approx 0.99$ close to the Thorne limit. Instead, GRMHD simulations of slim disks show that we may get a significantly lower equilibrium spin, $a_{*,\rm{max}}\approx 0.93$ (Gammie et al., 2004). This corresponds to $f_{\Lambda}^{\rm min}=0.95$. We adopted this value as a conservative assumption on the spin-up process and performed sensitivity checks. We find that our results are robust and practically unaffected for values even down to $f_{\Lambda}^{\rm min}=0.88$, corresponding to the extreme case of an equilibrium spin, $a_{*,\rm{max}}=0.90$, at $\dot{m}=10$ for every single accretion cycle at this rate. Even for $f_{\Lambda}^{\rm min}=0.80$ our derived spin-mass correlation, including the high-spin plateau, remains robust, while the saturation spin drops only by $\lesssim 4\%$.
@@
We note that if a magnetically arrested disk (MAD) (Bisnovatyi-Kogan and Ruzmaikin, 1974) forms, the accreting BH spins down due to angular momentum extraction via the Blandford-Znajek mechanism (Tchekhovskoy et al., 2011). Ricarte et al. (2023) have estimated that for thick disks the equilibrium spin drops down to $a_{*,{\rm max}}\approx 0.8,\,0.2$ for $\dot{m}=1,\,10$, respectively. However, in the dense stochastic environment of a gaseous proto-cluster the maintenance of a long-lived coherent MAD is uncertain. In addition, Ricarte et al. (2023) suggest that the spin equilibration timescale for MAD accretion attains the values $t_{\rm eq}^{\rm MAD}\approx 45,\,10,\,4.5\,{\rm Myr}$ for $\dot{m}=1,\,5,\,10$, respectively. Our mass-gap BHs have been accreting during their evolution at a rate of $\dot{m}\sim 1-3$ or less, while only the IMBH with $m_{\rm BH}\sim 10^{3}\,{\rm M}_{\odot}$ may have reached up to $\dot{m}\sim 10$. In our model the cluster gas is depleted within $13\,{\rm Myr}$; therefore, MAD accretion is certainly irrelevant for our mass-gap BHs ($150\,{\rm M}_{\odot}$) for which $t_{\rm eq}^{\rm MAD}>t_{\rm depletion}$. Most importantly, as we discuss in Appendix A, the BHs are disrupted at a timescale of $\Delta t_{\rm stoch}\approx 10^{-2}\,{\rm Myr}$. Therefore, both for mass-gap and IMBHs, even if a MAD disk forms, it cannot remain coherent after a few disruption events and does not have sufficient time during its lifetime to cause any significant spin-down. Overall, our model assumes that MAD accretion is negligible.
@@ 3.3 Spin-disk orientation
The accretion disk, if formed, gets erratically disrupted, repeatedly, every timestep, $\Delta t_{\rm stoch}$ (dynamically calculated as the system evolves), when a stochastic kick due to stellar encounters and gas turbulence is applied. Since the disk orientation depends on the BH motion, each disruption event drives a disk reorientation.
In order to evaluate the spin-disk relative orientation after every disruption event, we need the disk angular momentum, $J_{\rm d}\equiv J(R_{\rm d})$, where
@@
Therefore, we need a model for the surface density, $\Sigma$, and the disk radius, $R_{\rm d}$, which does not necessarily coincide with the circularization radius. We calculated the disk surface density, adopting the Shakura-Sunjaev model (Shakura and Sunyaev, 1973; Frank et al., 2002),
@@
where $f(R)=(1-(R/R_{\rm ISCO})^{-1/2})^{1/4}$ and the gravitational radius is $r_{\rm g}=Gm_{\rm BH}/c^{2}$. For thin disks we adopted the viscosity parameter $\alpha_{\rm thin}=0.1$. For slim disks, the Shakura-Sunjaev surface density can be a good approximation using an effective viscosity parameter of $\alpha_{\rm eff}=0.01$ (Sądowski et al., 2011a). Indeed, we verified that for $\alpha=0.01$, the Shakura-Sunyaev surface density matches the slim-disk GRRMHD simulations of Sądowski et al. (2011a) to within a factor of order unity. Therefore we adopted this as a sufficient analytic approximation. We also performed exploratory sensitivity checks over the range $\alpha_{\rm eff}=0.001-0.1$. The general trend of our derived spin-mass correlation is robust. A notable result is that higher $\alpha_{\rm eff}$ produces a somewhat broader BH spin distribution for the more massive BHs. Especially for $\alpha_{\rm eff}=0.1$ the IMBHs ($\sim 10^{3}{\rm M}_{\odot}$) exhibit reduced lower and upper spin boundaries and a nearly uniform spin distribution within this range: $a_{*}\sim 0.3-0.8$. For $m_{\rm BH}\lesssim 800\,{\rm M}_{\odot}$, however, there is no reduction in spin boundaries, but only a slight increase in their spread. Nevertheless, we emphasize that $\alpha_{\rm eff}=0.1$ is not strictly self-consistent within our slim-disk treatment, because it lowers the effective optical depth, thereby shifting the validity of the model below our super-Eddington regime for slim disks (Sądowski et al., 2011a).
@@
In order to define the disk radius, $R_{\rm d}$, we need first to introduce three characteristic radii. Firstly, we defined the outer radius, $R_{\rm out}$, as the maximum outer radius at which torques have enough time to be communicated, before the BH motion gets disrupted by a stochastic kick. It is defined by the condition $t_{\rm w}(R_{\rm out})=\Delta t_{\rm stoch}$, which gives
@@
$\frac{R_{\rm out}}{r_{\rm g}}=\left\{\begin{array}[]{ll}1.5\cdot 10^{8}\left(\frac{\Delta t_{\rm stoch}}{10^{-3}\,{\rm Myr}}\right)^{\frac{4}{5}}\left(\frac{m_{\rm BH}}{50{\rm M}_{\odot}}\right)^{-\frac{24}{25}}&,\;\mbox{thin}\\ 3.5\cdot 10^{8}\left(\frac{\Delta t_{\rm stoch}}{10^{-3}\,{\rm Myr}}\right)^{\frac{2}{3}}\left(\frac{H/R}{0.1}\right)^{\frac{2}{3}}\left(\frac{m_{\rm BH}}{50{\rm M}_{\odot}}\right)^{-\frac{2}{3}}&,\;\mbox{slim}\end{array}\right.$. (25)
@@
We define $t_{\rm w}$ and discuss timescales in Appendix B. It should be $R_{\rm d}\leq R_{\rm out}$; otherwise, there is not enough time for the torque to be communicated out to $R_{\rm d}$ within the stochastic timestep, $\Delta t_{\rm stoch}$.
@@
The Toomre radius is defined by the condition $Q(R_{\rm Q})=1$, where the Toomre parameter is $Q(R)=c_{\rm s,disk}\Omega_{\rm K,disk}(R)/(\pi G\Sigma(R))$. Stable disks are constrained by the condition $Q(R)>1$; that is, $R<R_{\rm Q}$. We have
@@
$\frac{R_{\rm Q}}{r_{\rm g}}=\left\{\begin{array}[]{ll}1.7\cdot 10^{10}\left(\frac{m_{\rm BH}}{50{\rm M}_{\odot}}\right)^{-\frac{52}{45}}\dot{m}^{-\frac{22}{45}}&,\;\mbox{thin}\\ 2.9\cdot 10^{9}\left(\frac{H/R}{0.1}\right)^{\frac{4}{5}}\left(\frac{m_{\rm BH}}{50{\rm M}_{\odot}}\right)^{-\frac{24}{25}}\dot{m}^{-\frac{14}{25}}&,\;\mbox{slim}\end{array}\right.$. (26)
It should also be $R_{\rm d}\leq R_{\rm Q}$; otherwise, the disk is unstable.
@@
At last, we defined the “breaking” radius $R_{\rm break}$. Numerical evidence suggests that the disk (either thin or slim) breaks into rings above a threshold radius (Nixon et al., 2012; Nealon et al., 2015),
@@
$\frac{R_{\rm break}}{r_{\rm g}}=\left(\frac{4a_{*}}{3\alpha(H/R)}\left|\sin\theta_{\rm rel}\right|\right)^{2/3}\,,\quad 45^{o}\lesssim\theta_{\rm rel}\lesssim 135^{o},$ (27)
@@
for such a severe misalignment (Nixon et al., 2012). The angle $\theta_{\rm rel}$ is the relative angle between the initial disk angular momentum (at the time a stochastic kick is applied) and BH-spin orientation,
$\cos\theta_{\rm rel}=-\hat{\bm{\omega}}_{\rm BH}\cdot\hat{\bm{J}}_{BH}|_{t=t_{\rm ini}},$ (28)
@@
since $\hat{\bm{J}}_{\rm disk}=-\hat{\bm{\omega}}_{\rm BH}$. It should be $R_{\rm d}\leq R_{\rm break}$ when $45^{o}\lesssim\theta_{\rm rel}\lesssim 135^{o}$; otherwise, the disk breaks into rings.
@@
Thus, $R_{\rm d}$ should be smaller than all four characteristic radii, $R_{\rm circ}$, $R_{\rm out}$, $R_{\rm Q}$, and $R_{\rm break}$; that is, we defined
$R_{\rm d}=\min\{R_{\rm circ},R_{\rm out},R_{\rm Q},R_{\rm break}\},$ (29)
@@
provided always that $R_{\rm ISCO}<R_{\rm circ}\leq R_{\rm acc}$, so that a disk is able to form. The $R_{\rm break}$ is considered only when $45^{o}\lesssim\theta_{\rm rel}\lesssim 135^{o}$; otherwise, it is ignored.
@@
For $R_{\rm acc}=R_{\rm B}$, which is the typical case for BHs close the cluster center, the circularization radius is